### v1

# Lagrangian multiforms and dispersionless integrable systems

Abstract
We demonstrate that interesting examples of Lagrangian multiforms appear naturally in the theory of multidimensional dispersionless integrable systems as (a) higher-order conservation laws of linearly degenerate PDEs in 3D, and (b) in the context of Gibbons-Tsarev equations governing hydrodynamic reductions of heavenly type equations in 4D.
MSC2020: 35Q51, 35Q75, 37K10, 37K25, 53C25, 53Z05.
Keywords: Lagrangian multiforms, higher conservation laws, dispersionless integrability, hydrodynamic reductions, Gibbons-Tsarev equations.
Department of Mathematical Sciences
Loughborough University
Loughborough, Leicestershire LE11 3TU
United Kingdom
E.V.Ferapontov@lboro.ac.uk
M.Vermeeren@lboro.ac.uk
1 Lagrangian multiform formalism
We consider Lagrangian multiforms
$\mathcal{L}=\sum_{i<j}L_{ij}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}.$
Here $L_{ij}$ are some functions of the jet variables $x^{i},x^{j},u,u_{i},u_{j},u_{ij}$ where $u$ is a function of the independent variables $x^{1},\dots,x^{n}$ and $u_{i}=u_{x^{i}},\ u_{ij}=u_{x^{i}x^{j}}$, etc, denote partial derivatives. We assume $L_{ji}=-L_{ij}$. For a surface $\Gamma$ in the space of independent variables, we consider the action integral $S_{\Gamma}=\int_{\Gamma}\mathcal{L}$. Lagrangian multiform theory requires that this action is critical for every choice of surface $\Gamma$, with respect to variations of $u$. (This requirement is also known as the pluri-Lagrangian principle.) This is the case if and only if the following three groups of multiform Euler-Lagrange equations hold:
$\frac{\delta_{ij}L_{ij}}{\delta u}=0,$ (1a)
$\frac{\delta_{ij}L_{ij}}{\delta u_{j}}-\frac{\delta_{ik}L_{ik}}{\delta u_{k}}=0,$ (1b)
$\frac{\delta_{ij}L_{ij}}{\delta u_{ij}}+\frac{\delta_{jk}L_{jk}}{\delta u_{jk}}+\frac{\delta_{ki}L_{ki}}{\delta u_{ki}}=0.$ (1c)
Here $\frac{\delta\cdots}{\delta\cdots}$ denote variational derivatives which, under the assumption that $L_{ij}$ does not depend on third or higher derivatives of $u$, can be computed as
$\frac{\delta_{ij}}{\delta u}=\frac{\partial}{\partial u}-\partial_{i}\frac{\partial}{\partial u_{i}}-\partial_{j}\frac{\partial}{\partial u_{j}}+\partial_{i}\partial_{j}\frac{\partial}{\partial u_{ij}},$
$\frac{\delta_{ij}}{\delta u_{j}}=\frac{\partial}{\partial u_{j}}-\partial_{i}\frac{\partial}{\partial u_{ij}},$
$\frac{\delta_{ij}}{\delta u_{ij}}=\frac{\partial}{\partial u_{ij}},$
where $\partial_{i}=\frac{\partial}{\partial x^{i}}$, etc.
If we allow the coefficients $L_{ij}$ to depend on derivatives with respect to $x^{k}$ for $k\neq i,j$ as well, but still assume no third or higher derivatives occur in $L_{ij}$, the system of multiform Euler-Lagrange equations (1a)–(1c) is extended by
$\frac{\delta_{ij}L_{ij}}{\delta u_{k}}$ $=0\qquad\forall k\neq i,j,$ (1d)
$\frac{\partial L_{ij}}{\partial u_{k\ell}}$ $=0\qquad\forall k,\ell\neq i,j,$ (1e)
$\frac{\partial L_{ij}}{\partial u_{j\ell}}-\frac{\partial L_{ik}}{\partial u_{k\ell}}$ $=0\qquad\forall\ell\neq i.$ (1f)
Note that the definition of variational derivatives remains the same: the additional derivatives are only taken in the variables corresponding to the indices of $\delta$. In particular:
$\frac{\delta_{ij}}{\delta u_{k}}=\frac{\partial}{\partial u_{k}}-\partial_{i}\frac{\partial}{\partial u_{ik}}-\partial_{j}\frac{\partial}{\partial u_{jk}}.$
For this framework to be nontrivial, we need to impose the condition that the multiform Euler-Lagrange equations (also known as multi-time Euler-Lagrange equations) are in involution, i.e. that the overdetermined system (1) can be solved, given suitable boundary conditions.
The fundamental idea of Lagrangian multiforms was established in [17]. The multiform Euler-Lagrange equations for a 2-form $\mathcal{L}$ depending on the second jet bundle first appeared in [31] and the multiform Euler-Lagrange equations for $\mathcal{L}$ depending on any jet bundle were derived in [32]. This derivation makes use of surfaces that are made up of flat pieces in coordinate directions. Then equation (1a) comes from the flat pieces, (1b) from the edges where two flat pieces meet, and (1c) from the corners where three or more flat pieces meet. For this reason, (1b) are called planar equations, (1b) edge equations and (1c) corner equations. Other derivations of the multiform Euler-Lagrange equations are possible. One can obtain equations (1) by restricting $\mathcal{L}$ to an arbitrary plane in the space of independent variables (not necessarily coordinate planes) and calculating the standard Euler-Lagrange equation, which should be satisfied for any such plane. Furthermore, in [30] it was pointed out that equivalent equations are obtained by taking pointwise variations of $\mathrm{d}\mathcal{L}$.
Lagrangian multiforms and conservation laws have been touched upon in [23, 29], where Lagrangian multiforms were constructed for systems with known variational symmetries. These papers use Noether’s theorem to obtain conservation laws from which a multiform is then constructed. In the present work, we take conservation laws as our starting point and investigate the additional structure provided by Lagrangian multiform theory.
In many examples, the exterior derivative of $\mathcal{L}$ factorises as
$\mathrm{d}\mathcal{L}=\sum_{i<j<k}A_{ijk}B_{ijk}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}\wedge\mathrm{d}x^{k},$ (2)
and it is observed that the system
$A_{ijk}=0,\quad B_{ijk}=0$ (3)
is equivalent to the full system of multiform Euler-Lagrange equations (1). For example, for the Lagrangian multiform
$\mathcal{L}=\sum_{i<j}(c^{i}-c^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}\ \mathrm{d}x^{i}\wedge\mathrm{d}x^{j},$ (4)
where $c^{i}=\text{const}$, the factorisation (2) holds with
$A_{ijk}$ $=(c^{i}-c^{j})\frac{u_{ij}}{u_{i}u_{j}}+(c^{j}-c^{k})\frac{u_{jk}}{u_{j}u_{k}}+(c^{k}-c^{i})\frac{u_{ik}}{u_{i}u_{k}},$
$B_{ijk}$ $=u_{ijk}-\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{i}k}\right),$
while the multiform Euler-Lagrange equations (1) are
$\!\left(\frac{u_{ij}^{2}}{u_{i}^{2}u_{j}}\right)_{i}+\left(\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}\right)_{j}+\left(\frac{2u_{ij}}{u_{i}u_{j}}\right)_{ij}=0,$ (5a)
$(c^{j}-c^{i})\left(\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}+2\frac{u_{ii}u_{ij}}{u_{i}^{2}u_{j}}-2\frac{u_{iij}}{u_{i}u_{j}}\right)-(c^{k}-c^{i})\left(\frac{u_{ik}^{2}}{u_{i}u_{k}^{2}}+2\frac{u_{ii}u_{ik}}{u_{i}^{2}u_{k}}-2\frac{u_{iik}}{u_{i}u_{k}}\right)=0,$ (5b)
$(c^{i}-c^{j})\frac{u_{ij}}{u_{i}u_{j}}+(c^{j}-c^{k})\frac{u_{jk}}{u_{j}u_{k}}+(c^{k}-c^{i})\frac{u_{ik}}{u_{i}u_{k}}=0.$ (5c)
This system is equivalent to the equations $A_{ijk}=0$, $B_{ijk}=0$ (a fact which is not immediately obvious, see Appendix A.1). Equations $A_{ijk}=0$ form the so-called Veronese web hierarchy [34],
$(c^{i}-c^{j})u_{k}u_{ij}+(c^{k}-c^{i})u_{j}u_{ik}+(c^{j}-c^{k})u_{i}u_{jk}=0,$ (6)
while equations $B_{ijk}=0$,
$u_{ijk}=\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right),$ (7)
characterise potential (Egorov) metrics $\sum_{i}u_{i}(\mathrm{d}x^{i})^{2}$ with diagonal curvature, meaning that all curvature components $R^{i}_{kkj}$ with distinct $i,j,k$ are identically zero. Equations (7) have appeared in [27] in the context of multidimensional consistency of partial differential equations. Remarkably, equations (6) and (7), although coming from entirely different geometric contexts, are in involution. This can be checked by computing the Mayer bracket of $A_{ijk}$ and $B_{ijk}$ [14]. Furthermore, the first set of Euler-Lagrange equations, (1a), coincide with the sigma-model governing harmonic maps of pseudo-Euclidean plane into a pseudo-Riemannian surface of constant curvature 1, see Appendix A.2.
Remark.
The Lagrangian multiform (4) possesses a translationally non-invariant version,
$\mathcal{L}=\sum_{i,j}(x^{i}-x^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}\ \mathrm{d}x^{i}\wedge\mathrm{d}x^{j},$ (8)
with the corresponding variational equations
$(x^{i}-x^{j})u_{k}u_{ij}+(x^{k}-x^{i})u_{j}u_{ik}+(x^{j}-x^{k})u_{i}u_{jk}=0,$ (9)
as well as (7), which remains unchanged. Note that translationally non-invariant Veronese web equation (9) has appeared in [15, 17].
That the system $A_{ijk}=0$, $B_{ijk}=0$ implies the multiform Euler-Lagrange equations can be proved in general. (The various elements of the proof below can be found in [23, 29, 30, 32].)
Theorem 1.
If the differential $\mathrm{d}\mathcal{L}$ factorises as in equation (2), then the full system of multiform Euler-Lagrange equations (1) follows from the system of equations $A_{ijk}=0,B_{ijk}=0.$
Proof.
Consider an arbitrary surface $\Gamma$ and a variation $v$ of $u$, leading to a variation of the action
$\delta S_{\Gamma}=\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\bigg{|}_{\varepsilon=0}\int_{\Gamma}\mathcal{L}[u+\varepsilon v].$
Without loss of generality we can assume that $v$ is supported on a small neighbourhood $\Omega\in\mathbb{R}^{n}$. Then $\delta S_{\Gamma}=\delta S_{\Gamma\cap\Omega}$ and we can find a three-dimensional volume $V$ such that inside of $\Omega$, the boundary of $V$ coincides with $\Gamma$, i.e. $\Gamma\cap\Omega=\partial V\cap\Omega$. Then
$\delta S_{\Gamma}$ $=\delta S_{\partial V\cap\Omega}=\delta S_{\partial V}=\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\bigg{|}_{\varepsilon=0}\int_{V}\mathrm{d}\mathcal{L}[u+\varepsilon v]$
$=\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\bigg{|}_{\varepsilon=0}\int_{V}\sum_{i<j<k}A_{ijk}[u+\varepsilon v]B_{ijk}[u+\varepsilon v]\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}\wedge\mathrm{d}x^{k}$
$=\int_{V}\sum_{i<j<k}\left(A_{ijk}[u]\frac{\mathrm{d}B_{ijk}[u+\varepsilon v]}{\mathrm{d}\varepsilon}+B_{ijk}[u]\frac{\mathrm{d}A_{ijk}[u+\varepsilon v]}{\mathrm{d}\varepsilon}\right)\!\bigg{|}_{\varepsilon=0}\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}\wedge\mathrm{d}x^{k},$
which is zero on solutions of the system $A_{ijk}=0,B_{ijk}=0$. Since $\Gamma$ was chosen arbitrarily, this implies that the multiform Euler-Lagrange equations hold. ∎
In the situation of Theorem 1, it is said that $\mathrm{d}\mathcal{L}$ attains a double zero on solutions of the system $A_{ijk}=0,\,B_{ijk}=0$. In other examples, the coefficients of $\mathrm{d}\mathcal{L}$ are sums of factorised terms,
$d\mathcal{L}=\sum_{i<j<k}\left(\sum_{\ell}A_{ijk}^{\ell}B_{ijk}^{\ell}\right)\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}\wedge\mathrm{d}x^{k},$
so that $\mathrm{d}\mathcal{L}$ attains a double zero on the system $A_{ijk}^{\ell}=0,\,B_{ijk}^{\ell}=0$. Again, this system implies the multiform Euler-Lagrange equations.
Factorisations of $\mathrm{d}\mathcal{L}$ were first used in a narrower sense in [23] and [29]. Since then, arguments involving double zeros have inspired the construction of Lagrangian multiforms in several different settings [28, 2, 19, 26].
The Lagrangian multiform (4) has two important properties, which motivate the structure of the paper:
  • The relation d=0(mod $\mathrm{d}\mathcal{L}=0\ (\text{mod }A_{ijk}=0)$ suggests that $\mathcal{L}$ is a (second-order) conservation law of the Veronese web hierarchy (6), for which $B_{ijk}$ can be viewed as its characteristic. In Section 2 we provide further examples of this kind by calculating second-order conservation laws of some linearly degenerate PDEs in 3D. It is yet to be explored to what extent conservation laws of integrable PDEs/hierarchies can be viewed as a common source of Lagrangian multiforms.
  • It will be shown in Section 3 that equations (6)-(7) coincide with Gibbons-Tsarev equations governing hydrodynamic reductions of the 4D second heavenly equation. This suggests that Gibbons-Tsarev equations in 4D constitute yet another natural source of Lagrangian multiforms.
2 Lagrangian multiforms and conservation laws of linearly degenerate PDEs
The examples below support a point of view of Lagrangian multiforms as (higher) conservation laws of integrable PDEs (hierarchies).
2.1 Veronese web hierarchy
The Veronese web hierarchy (6) possesses a second-order conservation law,
$\sum_{i<j}(c^{i}-c^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},$
as well as $n+1$ first-order conservation laws,
$\sum_{j\neq i}\frac{1}{c^{i}-c^{j}}\frac{u_{j}}{u_{i}}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},\qquad\sum_{i<j}(c^{i}-c^{j})u_{i}u_{j}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j};$
in the first of these expressions, the index $i\in\{1,\dots,n\}$ is fixed, while $j$ varies. Taking their linear combination gives Lagrangian multiform
$\mathcal{L}=\sum_{i<j}L_{ij}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},\qquad L_{ij}=(c^{i}-c^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}+\frac{1}{c^{i}-c^{j}}\left(n^{j}\frac{u_{i}}{u_{j}}+n^{i}\frac{u_{j}}{u_{i}}\right)+\varepsilon(c^{i}-c^{j})u_{i}u_{j},$
which can be seen as a multi-parameter deformation of the Lagrangian multiform (4). The exterior derivative factorises as in equation (2), with
$A_{ijk}=(c^{i}-c^{j})u_{k}u_{ij}+(c^{j}-c^{k})u_{i}u_{jk}+(c^{k}-c^{i})u_{j}u_{ik}=0$
and
$B_{ijk}$ $=u_{ijk}-\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right)$
$\quad+\frac{n^{i}u_{j}u_{k}}{2(c^{i}-c^{j})(c^{i}-c^{k})u_{i}}+\frac{n^{j}u_{i}u_{k}}{2(c^{j}-c^{i})(c^{j}-c^{k})u_{j}}+\frac{n^{k}u_{i}u_{j}}{2(c^{k}-c^{i})(c^{k}-c^{j})u_{k}}-\varepsilon\frac{u_{i}u_{j}u_{k}}{2}.$
Hence, by Theorem 1, the system $A_{ijk}=0$, $B_{ijk}=0$ implies the multiform Euler-Lagrange equations, which are (compare (5)):
$\!\left(\frac{u_{ij}^{2}}{u_{i}^{2}u_{j}}+\frac{1}{c^{i}-c^{j}}\left(\frac{n^{j}}{u_{j}}-\frac{n^{i}u_{j}}{u_{i}^{2}}\right)+\varepsilon(c^{i}-c^{j})u_{j}\right)_{i}$
$\qquad+\left(\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}+\frac{1}{c^{i}-c^{j}}\left(-\frac{n^{j}u_{i}}{u_{j}^{2}}+\frac{n^{i}}{u_{i}}\right)+\varepsilon(c^{i}-c^{j})u_{i}\right)_{j}+\left(\frac{2u_{ij}}{u_{i}u_{j}}\right)_{ij}=0,$
$(c^{j}-c^{i})\left(\frac{u_{ij}^{2}}{u_{i}u_{j}^{2}}+2\frac{u_{ii}u_{ij}}{u_{i}^{2}u_{j}}-2\frac{u_{iij}}{u_{i}u_{j}}-\frac{n^{i}}{u_{i}}-\frac{n^{j}u_{i}}{u_{j}^{2}}-\frac{n^{k}u_{i}}{u_{k}^{2}}+\varepsilon u_{i}\right)$
$\qquad-(c^{k}-c^{i})\left(\frac{u_{ik}^{2}}{u_{i}u_{k}^{2}}+2\frac{u_{ii}u_{ik}}{u_{i}^{2}u_{k}}-2\frac{u_{iik}}{u_{i}u_{k}}-\frac{n^{i}}{u_{i}}-\frac{n^{j}u_{i}}{u_{j}^{2}}-\frac{n^{k}u_{i}}{u_{k}^{2}}+\varepsilon u_{i}\right)=0,$
$(c^{i}-c^{j})\frac{u_{ij}}{u_{i}u_{j}}+(c^{j}-c^{k})\frac{u_{jk}}{u_{j}u_{k}}+(c^{k}-c^{i})\frac{u_{ik}}{u_{i}u_{k}}=0.$
2.2 Translationally non-invariant Veronese web hierarchy
The translationally non-invariant Veronese web hierarchy (9) also possesses a second-order conservation law,
$\sum_{i,j}(x^{i}-x^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},$
as well as $n+1$ first-order conservation laws,
$\sum_{j\neq i}\frac{1}{x^{i}-x^{j}}\frac{u_{j}}{u_{i}}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},\qquad\sum_{i,j}(x^{i}-x^{j})u_{i}u_{j}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j};$
in the first of these expressions, $i\in\{1,\dots,n\}$ is fixed, while $j$ varies. Taking their linear combination gives the Lagrangian multiform,
${\cal L}=\sum L_{ij}\,\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},\qquad L_{ij}=(x^{i}-x^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}+\frac{1}{x^{i}-x^{j}}\left(n^{j}\frac{u_{i}}{u_{j}}+n^{i}\frac{u_{j}}{u_{i}}\right)+\varepsilon(x^{i}-x^{j})u_{i}u_{j},$
which can be seen as a multi-parameter deformation of the Lagrangian multiform (8). This Lagrangian multiform, with $\varepsilon=0$, has appeared in [17]. The corresponding variational equations $A_{ijk}=0$ and $B_{ijk}=0$ are
$(x^{i}-x^{j})u_{k}u_{ij}+(x^{j}-x^{k})u_{i}u_{jk}+(x^{k}-x^{i})u_{j}u_{ik}=0$
and
$u_{ijk}=\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right)$
$-\frac{n^{i}u_{j}u_{k}}{2(x^{i}-x^{j})(x^{i}-x^{k})u_{i}}-\frac{n^{j}u_{i}u_{k}}{2(x^{j}-x^{i})(x^{j}-x^{k})u_{j}}-\frac{n^{k}u_{i}u_{j}}{2(x^{k}-x^{i})(x^{k}-x^{j})u_{k}}+\varepsilon\frac{u_{i}u_{j}u_{k}}{2},$
respectively.
2.3 Mikhalev equation
The Mikhalev equation has the form
$u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0.$ (10)
It first appeared in [18], in the context of the Hamiltonian formalism of KdV type hierarchies. It possesses three second-order conservation laws,
$\mathcal{L}=F\,\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}+G\,\mathrm{d}x^{3}\wedge\mathrm{d}x^{1}+H\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}.$
For each of them, the differential $\mathrm{d}\mathcal{L}=(F_{1}+G_{2}+H_{3})\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}$ factorises as
$F_{1}+G_{2}+H_{3}=(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13})\cdot(\Sigma),$
where the characteristic $\Sigma$ is a differential expression in $u$ such that the combined system $\{u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\ \Sigma=0\}$ is involutive. Thus, these conservation laws give rise to Lagrangian multiforms, where $L_{12}=H$, $L_{13}=-G$, and $L_{23}=F$.
Case 1.
Let
$F=u_{3}u_{11}^{2}-u_{13}^{2}-2u_{11}(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}),$
$G=-u_{11}^{2},$
$H=2u_{11}u_{13}-u_{1}u_{11}^{2}.$
Due to the factorisation
$F_{1}+G_{2}+H_{3}=-2(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13})\,u_{111},$
as well as the involutivity of the combined system,
$u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\qquad u_{111}=0,$ (11)
$\mathcal{L}$ has all properties of a Lagrangian multiform. The multiform Euler-Lagrange equations (1) in this case are all immediate consequences of the system (11). In particular, we find the two equations of (11) as multiform Euler-Lagrange equations of type (1e) and (1b):
$\frac{\delta_{23}F}{\delta u_{11}}=\frac{\partial F}{\partial u_{11}}=-2(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}),$
$\frac{\delta_{31}G}{\delta u_{1}}+\frac{\delta_{23}F}{\delta u_{2}}=-\partial_{1}\frac{\partial G}{\partial u_{11}}-0=2u_{111}.$
This example can be deformed by adding to $\mathcal{L}$ a first-order conservation law of the Mikhalev equation,
$\left(2u_{1}u_{3}^{2}-u_{1}^{3}u_{3}-u_{2}u_{3}\right)_{1}+\left(u_{1}^{3}-u_{1}u_{3}\right)_{2}+\left(u_{1}^{4}-3u_{1}^{2}u_{3}+u_{1}u_{2}+u_{3}^{2}\right)_{3}=0.$
Thus, we take
$\tilde{F}=u_{3}u_{11}^{2}-u_{13}^{2}-2u_{11}(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13})+2u_{1}u_{3}^{2}-u_{1}^{3}u_{3}-u_{2}u_{3},$
$\tilde{G}=-u_{11}^{2}+u_{1}^{3}-u_{1}u_{3},$
$\tilde{H}=2u_{11}u_{13}-u_{1}u_{11}^{2}+u_{1}^{4}-3u_{1}^{2}u_{3}+u_{1}u_{2}+u_{3}^{2},$
with the factorisation
$\tilde{F}_{1}+\tilde{G}_{2}+\tilde{H}_{3}=2\left(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}\right)\left(u_{3}-u_{111}-\frac{3}{2}u_{1}^{2}\right).$
Due to involutivity of the combined system,
$u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\qquad u_{3}-u_{111}-\frac{3}{2}u_{1}^{2}=0,$ (12)
the deformed ~$\tilde{\mathcal{L}}=\tilde{F}\,\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}+\tilde{G}\,\mathrm{d}x^{3}\wedge\mathrm{d}x^{1}+\tilde{H}\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}$ is also a Lagrangian multiform. Note that the second equation is the potential KdV equation.
Again, we find the two equations of (12) as multiform Euler-Lagrange equations of type (1e) and (1b):
$\frac{\delta_{23}\tilde{F}}{\delta u_{11}}=\frac{\partial\tilde{F}}{\partial u_{11}}=-2(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}),$
$\frac{\delta_{31}\tilde{G}}{\delta u_{1}}+\frac{\delta_{23}\tilde{F}}{\delta u_{2}}=\frac{\partial\tilde{G}}{\partial u_{1}}-\partial_{1}\frac{\partial\tilde{G}}{\partial u_{11}}+\frac{\partial\tilde{F}}{\partial u_{2}}=(3u_{1}^{2}-u_{3})+2u_{111}-u_{3},$
and all other multiform Euler-Lagrange equations are consequences of these.
Case 2.
Let
$F=(2u_{3}u_{11}-u_{12})u_{13}-u_{1}u_{3}u_{11}^{2}+2(u_{1}u_{11}-u_{13})(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}),$
$G=u_{1}u_{11}^{2}-u_{11}u_{13},$
$H=u_{1}^{2}u_{11}^{2}-2u_{1}u_{11}u_{13}-u_{3}u_{11}^{2}+u_{11}u_{12}+u_{13}^{2}.$
Due to the factorisation
$F_{1}+G_{2}+H_{3}=2\left(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}\right)\left(u_{1}u_{111}-u_{113}+\frac{1}{2}u_{11}^{2}\right),$
as well as the involutivity of the combined system,
$u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\qquad u_{1}u_{111}-u_{113}+\frac{1}{2}u_{11}^{2}=0,$
$\mathcal{L}$ is a Lagrangian multiform. Note that the second equation is equivalent to the Hunter-Saxton equation,
$(v_{3}-vv_{1})_{1}+\frac{1}{2}v_{1}^{2}=0,$
for $v=u_{1}$, which is an integrable PDE that arises in the theory of nematic liquid crystals [12]. This equation appears as a multiform Euler-Lagrange equation of type (1b):
$\frac{\delta_{13}G}{\delta u_{1}}-\frac{\delta_{23}F}{\delta u_{2}}=-u_{11}^{2}-2u_{1}u_{111}+2u_{113}=2\left((u_{13}-u_{1}u_{11})_{1}+\frac{1}{2}u_{11}^{2}\right).$
Case 3.
Let
$F=u_{3}(u_{3}-u_{1}^{2})u_{11}^{2}+2(u_{3}-u_{1}^{2})(u_{33}-u_{12}-u_{1}u_{13})u_{11}$
$\qquad-2u_{1}^{2}u_{13}^{2}+2u_{1}u_{13}u_{33}+u_{3}u_{13}^{2}+u_{12}(u_{12}-2u_{33}),$
$G=-(u_{1}u_{11}-u_{13})^{2},$
$H=(u_{1}u_{11}-u_{13})(-u_{1}^{2}u_{11}+u_{1}u_{13}+2u_{3}u_{11}-2u_{12}).$
Due to the factorisation
$F_{1}+G_{2}+H_{3}=-2(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13})(u_{112}+(u_{1}^{2}-u_{3})u_{111}-u_{1}u_{113}+u_{1}u_{11}^{2}-u_{11}u_{13}),$
as well as the involutivity of the combined system,
$u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\quad u_{112}+(u_{1}^{2}-u_{3})u_{111}-u_{1}u_{113}+u_{1}u_{11}^{2}-u_{11}u_{13}=0,$
$\mathcal{L}$ is a Lagrangian multiform. Note that, modulo (10), the second equation is equivalent to the Gurevich-Zybin equation,
$(\partial_{3}-v\,\partial_{1})^{2}v=0,$
for $v=u_{1}$, which is known to be linearisable by a reciprocal transformation, see [10, 11, 21]. This equation appears as a multiform Euler-Lagrange equation of type (1b):
$\frac{\delta_{13}G}{\delta u_{1}}-\frac{\delta_{23}F}{\delta u_{2}}=2u_{1}u_{11}^{2}+2u_{1}^{2}u_{111}-4u_{1}u_{113}-2u_{11}u_{13}+2u_{133}=2(\partial_{3}-u_{1}\,\partial_{1})^{2}u_{1}.$
An equivalent multiform can be obtained by subtracting $(u_{33}-u_{1}u_{13}+u_{11}u_{3}-u_{12})^{2}$ from $F$:
$\tilde{H}$ $=-u_{1}^{3}u_{11}^{2}+2u_{1}^{2}u_{11}u_{13}+2u_{1}u_{11}^{2}u_{3}-2u_{1}u_{11}u_{12}-u_{1}u_{13}^{2}-2u_{11}u_{13}u_{3}+2\,u_{12}u_{13},$
$\tilde{G}$ $=-\left(u_{1}^{2}u_{11}^{2}-2u_{1}u_{11}u_{13}+u_{13}^{2}\right),$
$\tilde{F}$ $=2u_{1}^{3}u_{11}u_{13}-u_{1}^{2}u_{11}^{2}u_{3}+2u_{1}^{2}u_{11}u_{12}-3u_{1}^{2}u_{13}^{2}-2u_{1}^{2}u_{11}u_{33}-2u_{1}u_{12}u_{13}+u_{13}^{2}u_{3}$
$\qquad+4u_{1}u_{13}u_{33}-u_{33}^{2}.$
Its exterior derivative factorises as
$\tilde{F}_{1}+\tilde{G}_{2}+\tilde{H}_{3}=-2(u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13})(u_{1}u_{11}^{2}+u_{1}^{2}u_{111}-2u_{1}u_{113}-u_{11}u_{13}+u_{133}),$
where the second factor is exactly the Gurevich-Zybin equation.
Similar second-order conservation laws, as well as the associated Lagrangian multiforms, can be constructed for other 3D linearly degenerate second-order dispersionless integrable PDE as classified in [6].
3 Lagrangian multiforms and Gibbons-Tsarev equations
In the most general setting, the method of hydrodynamic reductions (that leads to equations of Gibbons-Tsarev type) applies to quasilinear PDEs of the form
$\sum_{i=1}^{d}A^{i}(u)\,u_{x^{i}}=0,$ (13)
or any other classes of PDEs transformable into form (13), see below. Here $u$ is a (column) vector-function of $d$ independent variables $x^{1},\dots,x^{d}$, and $u_{x^{i}}$ denote partial derivatives (the matrices $A^{i}(u)$ do not need to be square). For definiteness, let us consider $4$-dimensional PDEs ($d=4$) with four independent variables $t,x,y,z$. Let us look for exact solutions of (13) in the form ${u}={u}(R)$ where the Riemann invariants $R=\{R^{1},...,R^{n}\}$ solve a triple of commuting diagonal systems
$R^{i}_{t}=\lambda^{i}(R)\,R^{i}_{x},\qquad R^{i}_{y}=\mu^{i}(R)\,R^{i}_{x},\qquad R^{i}_{z}=\eta^{i}(R)\,R^{i}_{x}.$ (14)
Note that the number $n$ of Riemann invariants is allowed to be arbitrary. Thus, the original multi-dimensional equation (13) is decoupled into a collection of commuting $(1+1)$-dimensional systems in Riemann invariants. Solutions of this type, known as nonlinear interactions of $n$ planar simple waves, were investigated in gas dynamics and magnetohydrodynamics [1, 22]. Later on, they reappeared in the context of the dispersionless KP hierarchy [13, 8, 9].
We recall, see [33], that the requirement of commutativity of the flows (14) is equivalent to the following restrictions on their characteristic speeds:
$\frac{\lambda_{j}^{i}}{\lambda^{j}-\lambda^{i}}=\frac{\mu_{j}^{i}}{\mu^{j}-\mu^{i}}=\frac{\eta_{j}^{i}}{\eta^{j}-\eta^{i}},$ (15)
$i\neq j,\ \lambda_{j}^{i}=\frac{\partial}{\partial{R^{j}}}\lambda^{i}$, etc. Substituting ${u}(R)$ into (13) and using (14), one arrives at an over-determined system of equations for $u(R)$ and the characteristic speeds $\lambda^{i}(R),\mu^{i}(R),\eta^{i}(R)$, known as Gibbons-Tsarev equations. These equations imply, in particular, that the characteristic speeds $\lambda^{i},\mu^{i}$ and $\eta^{i}$ must satisfy an algebraic relation which can be interpreted as the dispersion relation of system (13).
One can show that the maximum “amount” of $n$-component reductions a $d$-dimensional PDE may possess is parametrised, modulo changes of variables $R^{i}\to f^{i}(R^{i})$, by $(d-2)n$ arbitrary functions of one variable [3, 5].
A $d$-dimensional PDE is said to be integrable if its $n$-component reductions are locally parametrised by $(d-2)n$ arbitrary functions of one variable [3], [5]. Our main observation is that Gibbons-Tsarev type systems governing such reductions possess a Lagrangian multiform representation, in all dimensions $d\geq 4$. Below we provide Lagrangian multiform representations of Gibbons-Tsarev equations for some well-known integrable PDEs.
3.1 Second heavenly equation
The so-called Plebański second heavenly equation,
$\theta_{tx}+\theta_{zy}+\theta_{xx}\theta_{yy}-\theta^{2}_{xy}=0,$ (16)
describes self-dual Ricci-flat metrics of the form
$\mathrm{d}s^{2}=\mathrm{d}x\,\mathrm{d}t+\mathrm{d}y\,\mathrm{d}z+\theta_{yy}\,\mathrm{d}x^{2}-2\theta_{xy}\,\mathrm{d}x\,\mathrm{d}y+\theta_{xx}\,\mathrm{d}y^{2},$
see [24]. Hydrodynamic reductions of equation (16) were discussed in [3]. Introducing the notation $\theta_{xx}=u,$ $\theta_{xy}=v,$ $\theta_{yy}=w,$ $\theta_{tx}=p,$ $\theta_{zy}=v^{2}-uw-p$, one first rewrites (16) in quasilinear form (13),
$\begin{split}&u_{y}=v_{x},\qquad u_{t}=p_{x},\qquad v_{y}=w_{x},\qquad v_{t}=p_{y},\\ &v_{z}=(v^{2}-uw-p)_{x},\qquad w_{z}=(v^{2}-uw-p)_{y}.\end{split}$ (17)
Hydrodynamic reductions are sought in the form $u=u(R^{1},...,R^{n}),$ $v=v(R^{1},...,R^{n}),$ $w=w(R^{1},...,R^{n}),$ $p=p(R^{1},...,R^{n})$ where the Riemann invariants $R^{i}$ satisfy equations (14). For every solution of (14) we require that the corresponding $u,v,w,p$ solve (17). This implies
$p_{i}=\lambda^{i}u_{i},\qquad v_{i}=\mu^{i}u_{i},\qquad w_{i}=(\mu^{i})^{2}u_{i},$ (18)
along with the dispersion relation
$\lambda^{i}=2v\mu^{i}-w-u(\mu^{i})^{2}-\mu^{i}\eta^{i}.$ (19)
Recall that low indices in (18) indicate partial derivatives by the variables $R^{i}$. Substituting $\lambda^{i}$ into the commutativity conditions (15) and taking into account that the compatibility conditions for the relations $p_{i}=\lambda^{i}u_{i},\ v_{i}=\mu^{i}u_{i}$ imply
$u_{ij}=\frac{\mu_{j}^{i}}{\mu^{j}-\mu^{i}}u_{i}+\frac{\mu_{i}^{j}}{\mu^{i}-\mu^{j}}u_{j}.$
One arrives at the following Gibbons-Tsarev type equations:
$\begin{split}&\mu_{j}^{i}=\frac{(\mu^{j}-\mu^{i})^{2}}{\eta^{j}-\eta^{i}+u(\mu^{j}-\mu^{i})}\ u_{j},\\ &\eta^{i}_{j}=\frac{(\mu^{j}-\mu^{i})(\eta^{j}-\eta^{i})}{\eta^{j}-\eta^{i}+u(\mu^{j}-\mu^{i})}\ u_{j},\\ &u_{ij}=2\frac{\mu^{j}-\mu^{i}}{\eta^{j}-\eta^{i}+u(\mu^{j}-\mu^{i})}\ u_{i}u_{j}.\end{split}$ (20)
Solving equations (20) for $\mu^{i},\eta^{i}$ and $u$, determining $\lambda^{i}$ from (19) and calculating $p,v,w$ from equations (18) (which are automatically compatible by virtue of (20)), one obtains a general $n$-component hydrodynamic reduction of the second heavenly equation. Moreover, the commutativity conditions will also be satisfied identically.
We emphasize that system (20) is in involution and its general solution depends on $3n$ arbitrary functions of one variable. Indeed, one can arbitrarily prescribe the restrictions of $\mu^{i}$ and $\eta^{i}$ to the $R^{i}$-coordinate line. This gives $2n$ arbitrary functions. Moreover, one can arbitrarily prescribe the restriction of $u$ to each of the coordinate lines, which provides extra $n$ arbitrary functions. Since reparametrizations $R^{i}\to f^{i}(R^{i})$ leave the system (20) invariant, one concludes that general $n$-component reductions are locally parametrized by $2n$ arbitrary functions of one variable. This supports the evidence that the heavenly equation (16) is a four-dimensional integrable PDE [3].
Let us proceed with the analysis of the Gibbons-Tsarev system (20). Introducing $c^{i}=\eta^{i}+u\mu^{i}-v$, one readily obtains $c^{i}_{j}=0$ so that $c^{i}=c^{i}(R^{i})$ are arbitrary functions of the indicated variables. Ultimately, system (20) simplifies to
$\mu^{i}_{j}=\frac{(\mu^{j}-\mu^{i})^{2}}{c^{j}-c^{i}}\ u_{j},$ (21a)
$u_{ij}=2\frac{\mu^{j}-\mu^{i}}{c^{j}-c^{i}}\ u_{i}u_{j}.$ (21b)
The elimination of $\mu$’s from system (21) leads to equations for the variable $u$ alone. These can be obtained as follows. The equations (21b) lead to second-order PDEs for $u$ that form a translationally non-invariant deformation of the Veronese web hierarchy,
$(c^{j}-c^{i})u_{k}u_{ij}+(c^{i}-c^{k})u_{j}u_{ik}+(c^{k}-c^{j})u_{i}u_{jk}=0,$
see [34, 16, 17, 15]. Differentiating the equations (21b) by $R^{k}$, one obtains a collection of third-order PDEs (7) for $u$,
$u_{ijk}=\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right).$
Upon the identification $R^{i}\leftrightarrow x^{i}$, these equations correspond to Lagrangian multiform (4),
${\cal L}=\sum_{i,j}(c^{i}-c^{j})\frac{u_{ij}^{2}}{u_{i}u_{j}}\ \mathrm{d}R^{i}\wedge\mathrm{d}R^{j},$
where $c^{i}(R^{i})$ are arbitrary functions of the indicated variables (the choices $c^{i}=R^{i}$ and $c^{i}=const$ are of particular interest).
Lagrangian multiform for system (20).
System (20) is related to the Lagrangian multiform
$\mathcal{L}=\sum_{i,j}\left(\mu^{j}_{i}u_{j}-\mu^{i}_{j}u_{i}+\frac{(\mu^{j}-\mu^{i})^{2}}{c^{j}-c^{i}}u_{i}u_{j}\right)\mathrm{d}R^{i}\wedge\mathrm{d}R^{j}.$ (22)
Its multiform Euler-Lagrange equations are as follows:
  • (1b) yields
    $\mu_{j}^{i}+\frac{(\mu^{i}-\mu^{j})^{2}}{c^{i}-c^{j}}u_{j}=\mu_{j}^{k}+\frac{(\mu^{k}-\mu^{j})^{2}}{c^{k}-c^{j}}u_{j}.$
    In other words, this equation implies that the quantity
    $P_{j}:=\mu_{j}^{i}+\frac{(\mu^{i}-\mu^{j})^{2}}{c^{i}-c^{j}}u_{j}$
    does not depend on the choice of $i$.
  • (1a) yields
    $\frac{\partial P_{i}}{\partial R^{j}}=\frac{\partial P_{j}}{\partial R^{i}}.$
    It follows that there exists some function $P$ such that $P_{i}=\frac{\partial}{\partial R^{i}}P$ for all $i$.
  • (1c) is trivially satisfied: all three variational derivatives are identically zero.
  • (1a) with respect to $\mu^{i}$ instead of $u$ yields
    $-u_{ij}+2\frac{\mu^{j}-\mu^{i}}{c^{j}-c^{i}}u_{i}u_{j}=0.$
  • (1b) with respect to $\mu^{i}$ instead of $u$ yields the trivial equation $-u_{i}=-u_{i}$.
  • (1c) with respect to $\mu^{i}$ instead of $u$ is trivially satisfied: all three variational derivatives are identically zero.
Hence, the system of multiform Euler-Lagrange equations is equivalent to
$\mu_{j}^{i}=\frac{(\mu^{j}-\mu^{i})^{2}}{c^{j}-c^{i}}u_{i}+\frac{\partial P}{\partial R^{j}},$
$u_{ij}=2\frac{\mu^{j}-\mu^{i}}{c^{j}-c^{i}}u_{i}u_{j},$
which, except for the term $\frac{\partial P}{\partial R^{J}}$, matches equation (21). Note that we can remove this term by a gauge transformation $\tilde{\mu}^{i}=\mu^{i}-P$. For the Lagrangian multiform (22), each coefficients of $\mathrm{d}\mathcal{L}=\sum_{ijk}M_{ijk}\,\mathrm{d}R^{i}\wedge\mathrm{d}R^{j}\wedge\mathrm{d}R^{k}$ decomposes into a sum of three factorised terms:
$M_{ijk}$ $=\left(u_{ij}-2\frac{\mu^{j}-\mu^{i}}{c^{j}-c^{i}}u_{i}u_{j}\right)\left(\mu^{i}_{k}-\frac{(\mu^{k}-\mu^{i})^{2}}{c^{k}-c^{i}}u_{k}-\mu^{j}_{k}+\frac{(\mu^{k}-\mu^{j})^{2}}{c^{k}-c^{j}}u_{k}\right)$
$\qquad+\left(u_{jk}-2\frac{\mu^{k}-\mu^{j}}{c^{k}-c^{j}}u_{j}u_{k}\right)\left(\mu^{j}_{i}-\frac{(\mu^{i}-\mu^{j})^{2}}{c^{i}-c^{j}}u_{i}-\mu^{k}_{i}+\frac{(\mu^{i}-\mu^{k})^{2}}{c^{i}-c^{k}}u_{i}\right)$
$\qquad+\left(u_{ik}-2\frac{\mu^{i}-\mu^{k}}{c^{i}-c^{k}}u_{k}u_{i}\right)\left(\mu^{k}_{j}-\frac{(\mu^{j}-\mu^{k})^{2}}{c^{j}-c^{k}}u_{j}-\mu^{i}_{j}+\frac{(\mu^{j}-\mu^{i})^{2}}{c^{j}-c^{i}}u_{j}\right).$
3.2 First heavenly equation
The Plebański first heavenly equation,
$\Omega_{xy}\Omega_{zt}-\Omega_{xt}\Omega_{zy}=1,$ (23)
governs Kähler potentials of 4-dimensional self-dual Ricci-flat metrics
$\mathrm{d}s^{2}=2\Omega_{xy}\,\mathrm{d}x\,\mathrm{d}y+2\Omega_{zt}\,\mathrm{d}z\,\mathrm{d}t+2\Omega_{xt}\,\mathrm{d}x\,\mathrm{d}t+2\Omega_{zy}\,\mathrm{d}z\,\,\mathrm{d}y,$
see [24]. Hydrodynamic reductions of equation (23) were discussed in [7]. Using the notation $\Omega_{xy}=a,\ \Omega_{zt}=b,\ \Omega_{xt}=p,\ \Omega_{zy}=q$, equation (23) takes quasilinear form (13),
$a_{t}=p_{y},\qquad a_{z}=q_{x},\qquad b_{x}=p_{z},\qquad b_{y}=q_{t},\qquad ab-pq=1.$ (24)
Hydrodynamic reductions are sought in the form $a=a(R^{1},...,R^{n})$, $b=b(R^{1},...,R^{n})$, $p=p(R^{1},...,R^{n})$, $q=q(R^{1},...,R^{n})$, where the Riemann invariants $R^{i}$ solve a triple of commuting hydrodynamic type systems (14). The substitution into (24) implies
$q_{i}=\eta^{i}a_{i},\qquad p_{i}=\frac{\lambda^{i}}{\mu^{i}}a_{i},\qquad b_{i}=\frac{\eta^{i}\lambda^{i}}{\mu^{i}}a_{i}.$ (25)
Differentiation of $ab-pq=1$ implies $ab_{i}+ba_{i}=pq_{i}+qp_{i}$ which, by virtue of (25), gives the dispersion relation
$\lambda^{i}=\mu^{i}\frac{p\eta^{i}-b}{a\eta^{i}-q}.$ (26)
Substituting (26) back into (25), one obtains
$q_{i}=\eta^{i}a_{i},\qquad p_{i}=\frac{p\eta^{i}-b}{a\eta^{i}-q}a_{i},\qquad b_{i}=\eta^{i}\frac{p\eta^{i}-b}{a\eta^{i}-q}a_{i}.$ (27)
Calculation of the compatibility conditions for equations (27) results in
$\begin{split}&a_{ij}=\frac{\eta^{i}_{j}}{\eta^{j}-\eta^{i}}a_{i}+\frac{\eta^{j}_{i}}{\eta^{i}-\eta^{j}}a_{j},\\ &\eta^{i}_{j}(a\eta^{j}-q)a_{i}+\eta^{j}_{i}(a\eta^{i}-q)a_{j}=(\eta^{i}-\eta^{j})^{2}a_{i}a_{j}.\end{split}$ (28)
The commutativity conditions (15), with $\lambda^{i}$ given by (26), imply
$\eta^{i}_{j}=\frac{\mu^{i}(\eta^{j}-\eta^{i})^{2}}{\mu^{i}(a\eta^{j}-q)-\mu^{j}(a\eta^{i}-q)}a_{j}.$ (29)
Ultimately, combining (28), (15) and (29), we arrive at the following Gibbons-Tsarev type equations for $a,q,\mu^{i},\eta^{i}$:
$\begin{split}&q_{i}=\eta^{i}a_{i},\\ &a_{ij}=\frac{(\mu^{i}+\mu^{j})(\eta^{j}-\eta^{i})}{\mu^{i}(a\eta^{j}-q)-\mu^{j}(a\eta^{i}-q)}a_{i}a_{j},\\ &\eta^{i}_{j}=\frac{\mu^{i}(\eta^{j}-\eta^{i})^{2}}{\mu^{i}(a\eta^{j}-q)-\mu^{j}(a\eta^{i}-q)}a_{j},\\ &\mu^{i}_{j}=\frac{\mu^{i}(\eta^{j}-\eta^{i})(\mu^{j}-\mu^{i})}{\mu^{i}(a\eta^{j}-q)-\mu^{j}(a\eta^{i}-q)}a_{j},\end{split}$ (30)
where $i\neq j$. Note that $p,b,\lambda^{i}$ do not explicitly enter these equations. One can show by direct calculation that system (30) is in involution and its general solution depends, modulo reparametrisations $R^{i}\to f^{i}(R^{i})$, on $2n$ arbitrary functions of a single variable, thus confirming the integrability of the first heavenly equation.
Let us proceed with the analysis of the Gibbons-Tsarev system (30). Introducing potential $u$ by the formula $u_{i}=a_{i}/\mu^{i}$ (compatibility is guaranteed by (30)), as well as the functions $c^{i}=(a\eta^{i}-q)/\mu^{i}$, one can rewrite (30) in the form
$\begin{split}&q_{i}=\mu^{i}\eta^{i}u_{i},\\ &u_{ij}=2\frac{\eta^{j}-\eta^{i}}{c^{j}-c^{i}}\,u_{i}u_{j},\\ &\eta^{i}_{j}=\frac{(\eta^{j}-\eta^{i})^{2}}{c^{j}-c^{i}}\,u_{j},\\ &\mu^{i}_{j}=\frac{(\eta^{j}-\eta^{i})(\mu^{j}-\mu^{i})}{c^{j}-c^{i}}\,u_{j}.\end{split}$ (31)
It remains to verify the identity $c^{i}_{j}=0$, i.e. that the $c_{i}$ are functions of $R^{i}$ only, which makes the above equations for $u$ and $\eta^{i}$ identical to equations (21). Thus, both first and second heavenly equations correspond to one and the same Lagrangian multiform (which is not surprising as both equations describe one and the same class of self-dual Ricci-flat metrics, and are Bäcklund-related).
3.3 6D version of the second heavenly equation
A six-dimensional generalisation of the second heavenly equation,
$\theta_{t\tilde{t}}+\theta_{z\tilde{z}}+\theta_{tx}\theta_{zy}-\theta_{ty}\theta_{zx}=0,$ (32)
was proposed in [25]. Its hydrodynamic reductions were studied in [3]. Introducing the notation $\theta_{tx}=a,$ $\theta_{zy}=b,$ $\theta_{ty}=p,$ $\theta_{zx}=q,$ $\theta_{z\tilde{z}}=r,$ $\theta_{t\tilde{t}}=pq-ab-r$, one can rewrite (32) in quasilinear form (13),
$\begin{split}&a_{y}=p_{x},\qquad a_{z}=q_{t},\qquad b_{t}=p_{z},\qquad b_{x}=q_{y},\qquad b_{\tilde{z}}=r_{y},\qquad q_{\tilde{z}}=r_{x},\\ &p_{\tilde{z}}=(pq-ab-r)_{y}.\end{split}$ (33)
Hydrodynamic reductions are sought in the form $a=a(R^{1},...,R^{n}),$ $b=b(R^{1},...,R^{n}),$ $p=p(R^{1},...,R^{n}),$ $q=q(R^{1},...,R^{n}),$ $r=r(R^{1},...,R^{n})$ where the Riemann invariants $R^{1},...,R^{n}$ solve the commuting equations
$R^{i}_{x}=\lambda^{i}(R)\ R^{i}_{z},\qquad R^{i}_{y}=\mu^{i}(R)\ R^{i}_{z},\qquad R^{i}_{\tilde{z}}=\eta^{i}(R)\ R^{i}_{z},\qquad$
$R^{i}_{t}=\beta^{i}(R)\ R^{i}_{z},\qquad R^{i}_{\tilde{t}}=\gamma^{i}(R)\ R^{i}_{z}.$
The substitution into (33) implies
$\partial_{i}p=\beta^{i}\partial_{i}b,\qquad\partial_{i}r=\frac{\eta^{i}}{\mu^{i}}\partial_{i}b,\qquad\partial_{i}q=\frac{\lambda^{i}}{\mu^{i}}\partial_{i}b,\qquad\partial_{i}a=\frac{\lambda^{i}\beta^{i}}{\mu^{i}}\partial_{i}b,$ (34)
along with the dispersion relation
$\eta^{i}=\beta^{i}\mu^{i}q+\lambda^{i}p-\beta^{i}\lambda^{i}b-\mu^{i}a-\beta^{i}\gamma^{i}.$ (35)
Substituting $\eta^{i}$ into the commutativity conditions
$\frac{\partial_{j}\lambda^{i}}{\lambda^{j}-\lambda^{i}}=\frac{\partial_{j}\mu^{i}}{\mu^{j}-\mu^{i}}=\frac{\partial_{j}\eta^{i}}{\eta^{j}-\eta^{i}}=\frac{\partial_{j}\beta^{i}}{\beta^{j}-\beta^{i}}=\frac{\partial_{j}\gamma^{i}}{\gamma^{j}-\gamma^{i}},$
and taking into account that the compatibility conditions for the relations $\partial_{i}p=\beta^{i}\partial_{i}b$ imply
$\partial_{i}\partial_{j}b=\frac{\partial_{j}\beta^{i}}{\beta^{j}-\beta^{i}}\partial_{i}b+\frac{\partial_{i}\beta^{j}}{\beta^{i}-\beta^{j}}\partial_{j}b,$
one arrives at the following system:
$\begin{split}&\frac{\partial_{j}\beta^{i}}{\beta^{j}-\beta^{i}}=\frac{\partial_{j}\lambda^{i}}{\lambda^{j}-\lambda^{i}}=\frac{\partial_{j}\mu^{i}}{\mu^{j}-\mu^{i}}=\frac{\partial_{j}\gamma^{i}}{\gamma^{j}-\gamma^{i}}=\frac{\lambda^{i}-\lambda^{j}\mu^{i}/\mu^{j}}{q(\mu^{j}-\mu^{i})+b(\lambda^{i}-\lambda^{j})+\gamma^{i}-\gamma^{j}}\ \partial_{j}b,\\ &\partial_{i}\partial_{j}b=\frac{\lambda^{i}(1+\mu^{j}/\mu^{i})-\lambda^{j}(1+\mu^{i}/\mu^{j})}{q(\mu^{j}-\mu^{i})+b(\lambda^{i}-\lambda^{j})+\gamma^{i}-\gamma^{j}}\ \partial_{i}b\partial_{j}b.\end{split}$ (36)
Solving equations (36) for $\beta^{i}$, $\lambda^{i}$, $\mu^{i}$, $\gamma^{i}$ and $b$, determining $\eta^{i}$ from the dispersion relation (35) and calculating $p,r,q,a$ from the equations (34) (which are automatically compatible by virtue of (36)), one obtains a general $n$-component reduction of the equation (32). The commutativity conditions will be satisfied identically. System (36) is in involution and its general solution depends, up to reparametrisations $R^{i}\to\varphi^{i}(R^{i})$, on $4n$ arbitrary functions of one variable.
Let us proceed with the analysis of Gibbons-Tsarev system (36). Introducing the functions $c^{i}=\gamma^{i}+b\lambda^{i}-q\mu^{i}$, one can verify the identity $c^{i}_{j}=0$, thus, $c_{i}$ are functions of $R^{i}$ only, so that equations (36) take the form
$\begin{split}&\frac{\lambda^{i}_{j}}{\lambda^{j}-\lambda^{i}}=\frac{\mu^{i}_{j}}{\mu^{j}-\mu^{i}}=\frac{\lambda^{i}-\lambda^{j}\mu^{i}/\mu^{j}}{c^{i}-c^{j}}\,b_{j},\\ &b_{ij}=\frac{\lambda^{i}(1+\mu^{j}/\mu^{i})-\lambda^{j}(1+\mu^{i}/\mu^{j})}{c^{i}-c^{j}}\,b_{i}b_{j}.\end{split}$ (37)
Note that by introducing potential $u$ by the formula $u_{i}=\mu^{i}b_{i}$ (compatibility is guaranteed by (37)), as well as the functions $\phi^{i}=\lambda^{i}/\mu^{i}$, one can reduce (37) to the familiar form (21),
$\phi^{i}_{j}=\frac{(\phi^{j}-\phi^{i})^{2}}{c^{j}-c^{i}}\,u_{j},\qquad u_{ij}=2\frac{\phi^{j}-\phi^{i}}{c^{j}-c^{i}}\,u_{i}u_{j}.$
4 Concluding remarks
  • We considered (integrable) PDEs $F=0$ with a (higher) conservation law $C$ with non-constant characteristic (cosymmetry) $S$, so that $\mathrm{d}C=SF$. Our examples suggest that the constrained system, $F=S=0$, is often involutive. It would be interesting to clarify what additional conditions are required for this to be the case.
  • Equations (6) of the Veronese web hierarchy describe geometric objects known as Veronese webs, see [16, 4]. On the contrary, equations (7) describe potential (Egorov) metrics of diagonal curvature, typically arising in the theory of integrable systems of hydrodynamic type [33]. It seems remarkable that, put together, the equations governing these structures are compatible. In this connection, it would be interesting to understand the geometry of Veronese webs constrained by equations (7), as well as the properties of Egorov metrics constrained by equations (6).
  • Our examples suggest that Gibbons-Tsarev type systems governing hydrodynamic reductions of linearly degenerate dispersionless integrable PDEs possess a Lagrangian multiform representation. Such systems are only of interest in dimensions $d\geq 4$ as, according to [20], Gibbons-Tsarev systems of three-dimensional linearly degenerate PDEs are linearisable. In contrast, Gibbons-Tsarev systems of higher-dimensional linearly degenerate PDEs are not linearisable (not even Darboux integrable).
Acknowledgements
We thank S. Agafonov, V. Caudrelier, I. Krasilshchik, and M. Pavlov for useful discussions. MV is supported by the Engineering and Physical Sciences Research Council [Project reference EP/Y006712/1].
Appendix
A.1 Lagrangian multiform (4) and Euler-Lagrange equations
In this section we demonstrate by direct calculation that the multiform Euler-Lagrange equations (5) for the Lagrangian multiform (4) are equivalent to equations (6)–(7). The latter can be written as $A=0$ and $B=0$ with
$A=a^{j}{u_{ik}}{u_{j}}+a^{i}{u_{i}}{u_{jk}}+a^{k}{u_{ij}}{u_{k}},$
where $a^{i}=c^{j}-c^{k}$, etc, and
$B={u_{ijk}}-\frac{{u_{ij}}{u_{ik}}}{2{u_{i}}}-\frac{{u_{ij}}{u_{jk}}}{2{u_{j}}}-\frac{{u_{ik}}{u_{jk}}}{2{u_{k}}}.$
Note that $A=0$ is the corner equation (5c).
Lemma 2.
If $A=0$, then $B=0$ is equivalent to the edge equation (5b),
$\frac{1}{2}a^{k}{\left(\frac{u_{ij}^{2}}{{u_{i}}u_{j}^{2}}-\frac{2{u_{iij}}}{{u_{i}}{u_{j}}}+\frac{2{u_{ii}}{u_{ij}}}{u_{i}^{2}{u_{j}}}\right)}+\frac{1}{2}a^{j}{\left(\frac{{u_{ik}}^{2}}{{u_{i}}{u_{k}}^{2}}-\frac{2{u_{iik}}}{{u_{i}}{u_{k}}}+\frac{2{u_{ii}}{u_{ik}}}{u_{i}^{2}{u_{k}}}\right)}=0,$
and equations obtained from this by cyclic permutation of $(i,j,k)$.
Proof.
Taking into account that $a^{i}+a^{j}+a^{k}=0$, the left-hand side of equation (5b) can be written as
$\frac{1}{u_{i}u_{j}u_{k}}\left(A_{i}-a^{i}u_{i}B-\frac{1}{2}\frac{u_{ij}}{u_{j}}A-\frac{1}{2}\frac{u_{ik}}{u_{k}}A\right).\qed$
Lemma 3.
$A=0$ and $B=0$ together imply the planar equations (5a),
$a^{k}{\left(\frac{{u_{ij}}{u_{ijj}}}{{u_{i}}u_{j}^{2}}-\frac{{u_{iijj}}}{{u_{i}}{u_{j}}}+\frac{{u_{iij}}{u_{ij}}}{u_{i}^{2}{u_{j}}}-\frac{{u_{ii}}u_{ij}^{2}}{u_{i}^{3}{u_{j}}}+\frac{{u_{ii}}{u_{ijj}}}{u_{i}^{2}{u_{j}}}-\frac{u_{ij}^{2}{u_{jj}}}{{u_{i}}u_{j}^{3}}+\frac{{u_{iij}}{u_{jj}}}{{u_{i}}u_{j}^{2}}-\frac{{u_{ii}}{u_{ij}}{u_{jj}}}{u_{i}^{2}u_{j}^{2}}\right)}=0,$
and equations obtained from this by cyclic permutation of $(i,j,k)$.
Proof.
Differentiate $A=0$ with respect to $x^{i}$ and $x^{j}$ and simplify using $a^{i}+a^{j}+a^{k}=0$:
${\left({u_{i}}{u_{ijjk}}+{u_{ii}}{u_{jjk}}+{u_{iij}}{u_{jk}}\right)}a^{i}$ $+{\left({u_{ijj}}{u_{ik}}+{u_{iijk}}{u_{j}}+{u_{iik}}{u_{jj}}\right)}a^{j}$
$+{\left({u_{ijj}}{u_{ik}}+{u_{iij}}{u_{jk}}+{u_{iijj}}{u_{k}}\right)}a^{k}=0.$
We want to eliminate all derivatives with respect to $x^{k}$ from this equation. We can use $B=0$ and its differential consequences $B_{i}=0$ and $B_{j}=0$ to eliminate $u_{iijk}$ and $u_{ijjk}$, and (5b) to eliminate $u_{iik}$ and $u_{jjk}$. Then we simplify using $a^{i}+a^{j}+a^{k}=0$ and group terms strategically to obtain
$\frac{a^{i}}{2}{\left(\frac{{u_{iij}}}{u_{i}}+\frac{{u_{ijj}}}{{u_{j}}}+\frac{u_{ij}^{2}}{u_{i}{u_{j}}}+\frac{2{u_{ii}}{u_{jj}}}{u_{i}{u_{j}}}+\frac{{u_{ij}}{u_{ik}}}{u_{i}{u_{k}}}+\frac{{u_{ik}}{u_{jj}}}{{u_{j}}{u_{k}}}+\frac{{u_{ii}}{u_{jk}}}{u_{i}{u_{k}}}+\frac{{u_{ij}}{u_{jk}}}{{u_{j}}{u_{k}}}\right)}u_{i}u_{jk}$
$+\frac{a^{j}}{2}{\left(\frac{{u_{iij}}}{{u_{i}}}+\frac{u_{ijj}}{u_{j}}+\frac{u_{ij}^{2}}{{u_{i}}u_{j}}+\frac{2{u_{ii}}{u_{jj}}}{{u_{i}}u_{j}}+\frac{{u_{ij}}{u_{ik}}}{{u_{i}}{u_{k}}}+\frac{{u_{ik}}{u_{jj}}}{u_{j}{u_{k}}}+\frac{{u_{ii}}{u_{jk}}}{{u_{i}}{u_{k}}}+\frac{{u_{ij}}{u_{jk}}}{u_{j}{u_{k}}}\right)}u_{j}u_{ik}$
$+\frac{a^{k}}{2}{\left(-\frac{{u_{iij}}}{{u_{i}}}-\frac{{u_{ijj}}}{{u_{j}}}+\frac{u_{ij}^{2}}{{u_{i}}{u_{j}}}+\frac{2{u_{ii}}{u_{jj}}}{{u_{i}}{u_{j}}}+\frac{{u_{ij}}{u_{ik}}}{{u_{i}}u_{k}}+\frac{{u_{ik}}{u_{jj}}}{{u_{j}}u_{k}}+\frac{{u_{ii}}{u_{jk}}}{{u_{i}}u_{k}}+\frac{{u_{ij}}{u_{jk}}}{{u_{j}}u_{k}}\right)}u_{k}u_{ij}$
$+\frac{a^{k}}{2}{\left(2{u_{iijj}}-\frac{2{u_{ii}}{u_{ijj}}}{{u_{i}}}-\frac{2{u_{iij}}{u_{jj}}}{{u_{j}}}+\frac{2{u_{ii}}u_{ij}^{2}}{u_{i}^{2}}+\frac{2u_{ij}^{2}{u_{jj}}}{u_{j}^{2}}+\frac{2{u_{ii}}{u_{ij}}{u_{jj}}}{{u_{i}}{u_{j}}}\right)}u_{k}=0.$
Applying $A=0$ now allows us to eliminate $a^{i}$ and $a^{j}$:
$a^{k}\left(-\frac{{u_{iij}}u_{ij}}{{u_{i}}}-\frac{{u_{ijj}}u_{ij}}{{u_{j}}}\mathrel{+}{u_{iijj}}-\frac{{u_{ii}}{u_{ijj}}}{{u_{i}}}-\frac{{u_{iij}}{u_{jj}}}{{u_{j}}}+\frac{{u_{ii}}u_{ij}^{2}}{u_{i}^{2}}+\frac{u_{ij}^{2}{u_{jj}}}{u_{j}^{2}}+\frac{{u_{ii}}{u_{ij}}{u_{jj}}}{{u_{i}}{u_{j}}}\right)u_{k}=0.$
Finally, we divide by $-u_{i}u_{j}u_{k}$ to find the planar Euler-Lagrange equation (5a). ∎
Together, these Lemmas show that the system $A=0$, $B=0$ is equivalent to the system (5).
A.2 Lagrangian multiform (4) and sigma model equations
Here we provide some more detail on equation (5a), setting $(i,j)=(1,2)$:
$\left(\frac{u_{12}^{2}}{u_{1}^{2}u_{2}}\right)_{1}+\left(\frac{u_{12}^{2}}{u_{1}u_{2}^{2}}\right)_{2}+\left(\frac{2u_{12}}{u_{1}u_{2}}\right)_{12}=0.$
This fourth-order PDE can be rewritten as a second-order system,
$u_{12}-p\,u_{1}u_{2}=0,\qquad p_{12}+p\,p_{1}u_{2}+p\,p_{2}u_{1}+p^{3}u_{1}u_{2}=0,$
with the action functional
$\int\left(p^{2}u_{1}u_{2}+p_{1}u_{2}+p_{2}u_{1}\right)\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}.$
This Lagrangian governs harmonic maps from the pseudo-Euclidean ($x^{1},x^{2}$)-plane with the flat metric $\mathrm{d}x^{1}\,\mathrm{d}x^{2}$, to the pseudo-Riemannian manifold with coordinates $u,p$ and the metric $p^{2}\,\mathrm{d}u^{2}+2\,\mathrm{d}u\,\mathrm{d}p$. Note that the latter metric has constant curvature $1$; the nonzero Christoffel symbols are $\Gamma^{1}_{11}=-p,\ \Gamma^{2}_{12}=\Gamma^{2}_{21}=p,\ \Gamma^{2}_{11}=p^{3}$, where we label $u,p$ as the first and second coordinates, respectively.
In general, sigma-models describing harmonic maps from the pseudo-Euclidean ($x^{1},x^{2}$)-plane with the metric $\mathrm{d}x^{1}\,\mathrm{d}x^{2}$ to a pseudo-Riemannian manifold with coordinates $u^{i}$ and the metric $g_{ij}\,\mathrm{d}u^{i}\,\mathrm{d}u^{j}$, are governed by the Lagrangian $\int g_{ij}u^{i}_{1}u^{j}_{2}\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}$, with the Euler-Lagrange equations
$u^{i}_{12}+\Gamma^{i}_{jk}u^{j}_{1}u^{k}_{2}=0,$
where $\Gamma^{i}_{jk}$ are the Christoffel symbols of $g_{ij}$.